Find the distance of the origin from the plane which passes through (2, 1, 8) , (1, 0, 2) and (−3, 4, 6)


1
Expert's answer
2022-02-07T16:51:52-0500

For find the distance from the point to the plane we must find the distance from the point to the projection of the point onto a plane.

Lets find the equation of the plane. We are using this formula:



"\begin{vmatrix}
 x-x_1 & y-y_1 & z-z_1 \\x_2-x_1 & y_2-y_1 & z_2-z_1\\
 x_3-x_1 & y_3-y_1 & z_3-z_1
\end{vmatrix}=0"


where (x1, y1, z1), (x2, y2, z2) and (x3, y3, z3) are given points


"\begin{vmatrix}
 x-2 & y-1 & z-8 \\1-2 & 0-1 & 2-8\\
 -3-2 & 4-1 & 6-8
\end{vmatrix}=0"


"\begin{vmatrix}
 x-2 & y-1 & z-8 \\-1 & -1 & -6\\
 -5 & 3 & -2
\end{vmatrix}=0"

"(x-2)((-1)(-2)-(-6)3)-(y-1)((-1)(-2)-(-6)(-5)) + (z -8)((-1)3-(-1)(-5)) = 0"


"20(x-2)+28(y-1)+(-8)(z-8)=0"

"20x+28y-8z-4=0"

After simplifying equation of the plane is "5x+7y-2z-1=0"


Distance of the origin O(0, 0, 0) from the plane


"Ax+By+Cz+D=0" could be find according to formula


"d=\frac{|A\cdot 0+B\cdot 0+C\cdot 0+D|}{\sqrt{A^2+B^2+C^2}}"

In our case



"d=\frac{|5\cdot 0+7\cdot 0-2\cdot 0-1|}{\sqrt{5^2+7^2+2^2}}=\frac{1}{\sqrt{78}}\approx 0.113"

Answer: the distance of the origin from the plane which passes through (2, 1, 8) , (1, 0, 2) and (−3, 4, 6) is equal "\frac{1}{\sqrt{78}}\approx 0.113" .